# A Time-Scaled ETAS Model for Earthquake Forecasting

### v1

@@ Abstract
The Himalayan region, particularly Nepal, is highly susceptible to frequent and severe seismic activity, underscoring the urgent need for robust earthquake forecasting models. This study introduces a suite of time-scaled Epidemic-Type Aftershock Sequence (ETAS) models tailored for earthquake forecasting in Nepal, leveraging seismic data from 2000 to 2020. By incorporating alternative time-scaling approaches—such as calibration, proportional hazards, log-linear, and power time scales—the models capture nuanced temporal patterns of aftershocks, improving event classification between background and triggered occurrences. We evaluate model performance under various assumptions of earthquake magnitude distributions (exponential, gamma, and radially symmetric) and employ optimization techniques including the Davidon-Fletcher-Powell algorithm and Iterative Stochastic De-clustering. The results reveal that time-scaling significantly enhances model interpretability and predictive accuracy, with the ISDM-based ETAS model achieving the best fit. This work not only deepens the statistical understanding of earthquake dynamics in Nepal but also lays a foundation for implementing more effective early warning systems in seismically active regions.

### v2

@@ Abstract
The Himalayan region, including Nepal, is prone to frequent and large earthquakes. Accurate forecasting of these earthquakes is crucial for minimizing loss of life and damage to infrastructure. In this study, we propose various time-scaled Epidemic Type Aftershock Sequence (ETAS) models to forecast earthquakes in Nepal. The ETAS model is a statistical model that describes the temporal and spatial patterns of aftershocks following a main shock. A dataset of earthquake occurrences in Nepal from 2000 to 2020 was collected, and this data was used to fit the models showcased in this article. Our results show that the time-scaled ETAS model is able to accurately forecast earthquake occurrences in Nepal, and could be a useful tool for earthquake early warning systems in the region.